### v1

@@ 4 Monte-Carlo Simulation Results
We consider three estimators: the simple DTE (Simple) and the DTE with regression adjustment based on the linear regression (OLS) and logit model (Logit). To evaluate the finite sample performance of these DTE estimators, we compute their bias and root mean squared error (RMSE) across 10,000 simulations. Additionally, we assess the average length and coverage probability of their 95% pointwise confidence intervals. In the main text, we present results for DGP1 and DGP3 under a treatment assignment probability of $\pi_{1}=0.5$. Figures displaying results for all other combinations of DGPs and treatment assignment probabilities can be found in Supplemental Material B.
@@ 5.1 Nudge and Water Consumption
Figure 3 presents the DTE and PTE of the intervention in comparison to the control group. We compute the DTE for $y\in\{0,1,2,\dots,200\}$. The top left panel of Figure 3 displays the simple estimate of the DTE, whereas the top right panel illustrates the regression-adjusted estimate of the DTE obtained with logit model. The shaded areas denote the 95% pointwise confidence intervals. The results indicate that the regression-adjusted DTE has around 4% to 32% smaller standard errors and consequently tighter confidence intervals compared to the simple DTE.
@@ A.1 Proof of Theorem 1
$\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}=\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-2\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(1)}^{simple},(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\Big{)}$
@@
Since $G_{Y(k)|\bm{X}}$ belongs to the canonical class, we have that $\mathbb{E}[G_{Y(k)|\bm{X}}]=F_{Y(k)}$ and thus we can show that $\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}=\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}^{2}\big{]}$. Here, the equality hold only when $F_{Y(k)|\bm{X}}=F_{Y(k)}$ or $\bm{X}$ has no predictive power for the event $1\!{\rm l}_{\{Y(k)\leq\cdot\}}$.
@@ A.2 Distributional Regression
with $\Lambda(\cdot)$ being the link function and $R(\cdot)=\lambda(\cdot)/\{\Lambda(\cdot)[1-\Lambda(\cdot)]\}$. The class $\mathcal{H}$ consists of a Lipschitz transformation of VC-subgraph classes with Lipschitz coefficients and bounded from above by $\|T(\bm{X})\|$ up to some constant factor. Also, an envelop function for $\mathcal{H}$ is bounded from above by $\|T(\bm{X})\|$ up to some constant factor and thus is square integrable under Assumption A1. A Lipschitz composition of a Donsker class is a Donsker class. See 19.20 of van der Vaart (2000). Thus, the desired result for the first class is obtained. A simpler argument prove the desired conclusion for the other class. ∎
@@ A.3 Proof of Theorem 2
For treatment $k\in\mathcal{K}$, we have that $\widehat{\mathbb{F}}_{Y(k)}^{simple}=\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\widehat{G}_{Y(k)|\bm{X}}$, provided that the link function is of the canonical class, and we can write
@@
$\sqrt{n}\left(\begin{array}[]{c}\widehat{F}_{Y(1)}-F_{Y(1)}\\ \vdots\\ \widehat{F}_{Y(K)}-F_{Y(K)}\end{array}\right)\rightsquigarrow\mathbb{B}:=\left[\begin{array}[]{c}\mathbb{B}_{1}\\ \vdots\\ \mathbb{B}_{K}\\ \end{array}\right]\ \ \mathrm{in}\ \ell^{\infty}(\mathcal{X}{\times}\mathcal{Y})^{2},$
@@
As $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear transformations, the conclusion follows. ∎
@@ A.4 Proof of Theorem 3
Applying a similar argument in the proof of Theorem 2, we can show that
@@
in $\ell^{\infty}(\mathcal{Y})$. Furthermore, the transformations $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear in the regression-adjusted estimator, the desired conclusion follows. ∎
@@ Appendix B Additional Simulation Results
Note: (DGP1, continuous outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
Note: (DGP2, continuous outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
Note: (DGP2, continuous outcome, $\pi_{1}=0.5$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
Note: (DGP3, discrete outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
Note: (DGP4, discrete outcome, $\pi_{1}=0.3$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
Note: (DGP4, discrete outcome, $\pi_{1}=0.5$): Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.

### v2

@@ 4 Monte-Carlo Simulation Results
We consider three estimators: the simple DTE (Simple) and the DTE with regression adjustment based on the linear regression (OLS) and logit model (Logit). To evaluate the finite sample performance of these DTE estimators, we compute their bias and root mean squared error (RMSE) across 10,000 simulations. Additionally, we assess the average length and coverage probability of their 95% pointwise confidence intervals, based on analytic standard errors derived from the sample analogs. In the main text, we present results for DGP1 and DGP3 under a treatment assignment probability of $\pi_{1}=0.5$. Figures displaying results for all other combinations of DGPs and treatment assignment probabilities can be found in Supplemental Material B. Additional results using bootstrap confidence intervals and polynomial transformations of covariates are presented in Sections B.1 and B.2, respectively.
@@ 5.1 Nudge and Water Consumption
Figure 3 presents the DTE and PTE of the intervention in comparison to the control group. We compute the DTE for $y\in\{0,1,2,\dots,200\}$. The top left panel of Figure 3 displays the simple estimate of the DTE, whereas the top right panel illustrates the regression-adjusted estimate of the DTE obtained with logit model. The shaded areas denote the 95% pointwise confidence intervals, based on analytic standard errors derived from the sample analogs. The results indicate that the regression-adjusted DTE has around 4% to 32% smaller standard errors and consequently tighter confidence intervals compared to the simple DTE.
@@ A.1 Proof of Theorem 1
$\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}=\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}-2\mathrm{Cov}\Big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple},(\widehat{\mathbb{P}}_{\bm{X}}^{(k)}-\widehat{\mathbb{P}}_{\bm{X}})G_{Y(k)|\bm{X}}\Big{)}$
@@
Under the assumption that GLMs use a canonical link function, the conditional distribution is an unbiased estimator of the unconditional distribution or $\mathbb{E}[G_{Y(k)|\bm{X}}]=F_{Y(k)}$, as explained in Section A.5. Thus we can show that $\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}=\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}^{2}\big{]}$. Also, we have that $\mathrm{Var}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}\big{)}=\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-F_{Y(k)}\big{)}^{2}\big{]}$. Combining these results with the mathematical display above leads to that
$\mathrm{Var}\big{(}\widehat{\mathbb{F}}_{Y(k)}^{simple}\big{)}\geq\mathrm{Var}\big{(}\widetilde{\mathbb{F}}_{Y(k)}\big{)}+o(n^{-1}),$
provided that $\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-F_{Y(k)}\big{)}^{2}\big{]}\geq\mathbb{E}\big{[}\big{(}1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}}\big{)}^{2}\big{]}$.
@@ A.2 Distributional Regression
with $\Lambda(\cdot)$ being the inverse link function and $R(\cdot)=\lambda(\cdot)/\{\Lambda(\cdot)[1-\Lambda(\cdot)]\}$. The class $\mathcal{H}$ consists of a Lipschitz transformation of VC-subgraph classes with Lipschitz coefficients and bounded from above by $\|T(\bm{X})\|$ up to some constant factor. Also, an envelop function for $\mathcal{H}$ is bounded from above by $\|T(\bm{X})\|$ up to some constant factor and thus is square integrable under Assumption A1. A Lipschitz composition of a Donsker class is a Donsker class. See 19.20 of van der Vaart (2000). Thus, the desired result for the first class is obtained. A simpler argument prove the desired conclusion for the other class. ∎
@@ A.3 Proof of Theorem 2
(a) For treatment $k\in\mathcal{K}$, we have that $\widehat{\mathbb{F}}_{Y(k)}^{simple}=\widehat{\mathbb{P}}_{\bm{X}}^{(k)}\widehat{G}_{Y(k)|\bm{X}}$, provided that the link function is a canonical link function, and we can write
@@
$\sqrt{n}\left(\begin{array}[]{c}\widehat{F}_{Y(1)}-F_{Y(1)}\\ \vdots\\ \widehat{F}_{Y(K)}-F_{Y(K)}\end{array}\right)\rightsquigarrow\mathbb{B}:=\left[\begin{array}[]{c}\mathbb{B}_{1}\\ \vdots\\ \mathbb{B}_{K}\\ \end{array}\right]\ \ \mathrm{in}\ \ell^{\infty}(\mathcal{X}{\times}\mathcal{Y})^{2},$ (A.18)
@@
As $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear transformations, the conclusion follows.
(b) The QTE is obtained through a difference between two inverse maps: $\Psi^{QTE}_{k,k^{\prime}}(\gamma)=F{Y_{k}}^{-1}(\gamma)-F_{Y_{k^{\prime}}}^{-1}(\gamma)$ for $\gamma:=(F_{Y_{1}},\dots,F_{Y_{K}})^{\top}$. Given that $F_{Y_{k}}$ and $F_{Y_{k^{\prime}}}$ are continuously differentiable with strictly positive densities $f_{Y_{k}}$ and $f_{Y_{k^{\prime}}}$ on their respective supports, we can show that the functional map $\Psi^{QTE}_{k,k^{\prime}}$ is Hadamard differentiable, and its Hadamard derivative $\psi^{QTE}_{k,k^{\prime},\gamma}$ at $\gamma$ is defined in the statement of this theorem (see Lemma 3.9.20, van der Vaart and Wellner, 1996). Applying the function delta method (Theorem 3.9.4, van der Vaart and Wellner, 1996), we can obtain the desired result. ∎
@@ A.4 Proof of Theorem 3
(a) Applying a similar argument in the proof of Theorem 2, we can show that
@@
in $\ell^{\infty}(\mathcal{Y})$. Furthermore, the transformations $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ and $\Psi^{PTE}_{k,k^{\prime},h}(\cdot)$ are linear in the regression-adjusted estimator, the desired conclusion follows.
(b) The desired result follows by combining the argument presented in the proof of Theorem 2 with the result established in part (a) above. ∎
A.5 Generalized Linear Models and Canonical Link Functions
This subsection provides a brief overview of generalized linear models (GLMs) with particular emphasis on the canonical link function. We begin with the fundamental structure of GLMs, discuss their connection to exponential families, and then focus on the properties of canonical link functions. To maintain generality, we use notation distinct from that in the rest of this paper. For a comprehensive treatment of GLMs, readers are referred to McCullagh and Nelder (1989) among others.
We consider a scalar random variable $Y$ and a vector of covariates $X$. The essential structure of GLM for modeling the relationship between $Y$ and $X$ is
$g(\mathbb{E}[Y|X])=X^{\top}\beta,$
where $g:\mathbb{R}\to\mathbb{R}$ is the link function and $\beta$ is a $k\times 1$ vector of unknown parameters. Note that the linear regression model $\mathbb{E}[Y|X]=X^{\top}\beta$ is a special case where $g(\cdot)$ is the identity map $g(\mu)=\mu$.
GLMs are fundamentally connected to exponential family distributions, which encompass commonly-used probability distributions through a unified representation of their density:
$f(y;\theta,\phi)=\exp\left(\frac{y\theta-b(\theta)}{a(\phi)}+c(y,\phi)\right).$
Here, $\theta$ is the canonical parameter and captures the essential relationship between the response and its mean. The function $b(\theta)$ is the cumulant function, taking specific forms for different distributions in the family. The dispersion parameter $\phi$ enters through the scale function $a(\phi)$, while $c(y,\phi)$ ensures proper normalization of the density. Under the exponential family, we have
$\mathbb{E}[Y|X]=b^{\prime}(\theta)\ \ \mathrm{and}\ \ Var(Y|X)=b^{\prime\prime}(\theta)a(\phi),$
where $b^{\prime}(\cdot)$ and $b^{\prime\prime}(\cdot)$ denote the first and second derivatives of $b(\cdot)$, respectively.
There are various choices for the link function $g(\cdot)$ in GLMs, with the canonical link function being a popular choice that relates the linear predictor $X^{\top}\beta$ to the canonical parameter $\theta$. Specifically, let $(b^{\prime})^{-1}:\mathbb{R}\to\mathbb{R}$ be the inverse function of $b^{\prime}(\cdot)$ and then the math display above imples that $(b^{\prime})^{-1}(\mathbb{E}[Y|X])=\theta$ under the exponential family. The link function $g(\cdot)$ is called canonical when $g(\cdot)=(b^{\prime})^{-1}(\cdot)$. This canonical link function yields the relationship between the canonical parameter and the linear predictor: $\theta=X^{\top}\beta$. The form of the canonical link varies across different distributions in the exponential family. For the normal distribution, it is the identity function $g(\mu)=\mu$, which yields the standard linear regression model. For the Bernoulli distribution, it is the logit function $g(\mu)=\log(\mu/(1-\mu))$, mapping from the probability space $(0,1)$ to the real line. For the Poisson distribution, it is the logarithm function $g(\mu)=\log(\mu)$, which ensures the predicted mean remains positive.
We consider a random sample $\{(Y_{i},X_{i})\in\mathbb{R}\times\mathbb{R}^{k}\}_{i=1}^{n}$ with sample size $n$ as independent copies of $(Y,X)$. Under the canonical link function, the log-likelihood is
$\ell(\beta,\phi):=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{Y_{i}(X_{i}^{\top}\beta)-b(X_{i}^{\top}\beta)}{a(\phi)}+c(Y_{i},\phi)\right).$
The canonical link function ensures uniqueness of the maximum likelihood estimates (when they exist) and yields simpler expressions for the score statistic and information matrix. A key property of the canonical link function is its relationship to the score function. The score function for $\beta$ is
$\frac{\partial\ell(\beta,\phi)}{\partial\beta}$ $=\frac{1}{a(\phi)}\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-b^{\prime}(X_{i}^{\top}\beta)\big{)}X_{i}$
$=\frac{1}{a(\phi)}\frac{1}{n}\sum_{i=1}^{n}\big{(}Y_{i}-g^{-1}(X_{i}^{\top}\beta)\big{)}X_{i}.$
When $X$ includes a constant term, the maximum likelihood estimator $\hat{\beta}$ satisfying $\partial\ell(\hat{\beta},\hat{\phi})/\partial\beta=0$ implies $n^{-1}\sum_{i=1}^{n}\big{(}Y_{i}-g^{-1}(X_{i}^{\top}\hat{\beta})\big{)}=0$ and thus
$\frac{1}{n}\sum_{i=1}^{n}Y_{i}=\frac{1}{n}\sum_{i=1}^{n}g^{-1}(X_{i}^{\top}\hat{\beta}).$
Consequently, the canonical link function leads to an unbiased estimator of the unconditional mean.
A.6 Semiparametric Efficiency Bound
In this section, we derive the semiparametric efficiency bound of the DTE at a given $y\in\mathcal{Y}$, denoted by $\Delta^{DTE}_{k,k^{\prime}}(y)$. [^14] We thank an anonymous referee for raising this interesting question. The approach is analogous to that of Hahn (1998).
First, we characterize the tangent space. To that end, the joint density of the observed variables $(Y,\bm{W},\bm{X})$ can be written as:
$f(y,w,x)=f(y|w,x)f(w|x)f(x)=\prod_{k=1}^{K}\{f_{k}(y|x)\pi_{k}\}^{W_{k}}f(x),$
where $f_{k}(y|x):=P(Y=y|W_{k}=1,X=x)$ and $\pi_{k}=P(W_{k}=1|X=x)$ for all $x\in\mathcal{X}$. Notice that $f(w|x)$ does not depend on $x$ in our setup.
Consider a regular parametric submodel indexed by $\theta$:
$f(y,w,x;\theta):=\prod_{k=1}^{K}\{f_{k}(y|x;\theta)\pi_{k}(\theta)\}^{W_{k}}f(x;\theta),$
which equals $f(y,w,x)$ when $\theta=\theta_{0}$.
The corresponding score of $f(y,w,x;\theta)$ is given by
$s(y,w,x;\theta):=$ $\frac{\partial\ln f(y,w,x;\theta)}{\partial\theta}$
$=$ $\sum_{k=1}^{K}\Big{(}W_{k}\dot{f}_{k}(y|x;\theta)+W_{k}\dot{\pi}_{k}(\theta)\Big{)}+\dot{f}(x;\theta),$
where $\dot{f}$ denotes a derivative of the log, i.e., $\dot{f}_{k}(y|x;\theta)=\frac{\partial\ln f_{k}(y|x;\theta)}{\partial\theta}$, $\dot{\pi}_{k}(\theta)=\frac{\partial\ln\pi_{k}(\theta)}{\partial\theta}$ and $\dot{f}(x;\theta)=\frac{\partial\ln f(x;\theta)}{\partial\theta}$.
At the true value, the expectation of the score equals zero. The tangent space of the model is the set of functions that are mean zero and satisfy the additive structure of the score:
$\mathcal{T}=\Big{\{}\sum_{k=1}^{K}\Big{(}W_{k}s_{k}(y|x)+W_{k}s_{\pi}\Big{)}+s_{x}(x)\Big{\}},$ (A.19)
where $s_{k}(y|x)$ and $s_{x}(x)$ are mean-zero functions and $s_{\pi}$ is a real number.
The semiparametric variance bound of $\Delta^{DTE}_{k,k^{\prime}}(y)$ is the variance of the projection onto the tangent space $\mathcal{T}$ of a mean-zero function $\psi(Y,\bm{W},\bm{X})$ with finite second order moments, which satisfies for all regular parametric submodels
$\frac{\partial\Delta^{DTE}_{k,k^{\prime}}(y;F_{\theta})}{\partial\theta}\Big{|}_{\theta=\theta_{0}}=E[\psi(Y,\bm{W},\bm{X})\cdot s(Y,\bm{W},\bm{X})]\Big{|}_{\theta=\theta_{0}}.$ (A.20)
If $\psi$ itself already lies in the tangent space, the variance bound is given by $E[\psi^{2}].$ Now, the DTE is
[…]
Letting $\gamma_{k}^{y}(\bm{X}):=E[1\!{\rm l}_{\{Y(k)\leq y\}}|\bm{X}]$ and $\gamma_{k,k^{\prime}}^{y}(\bm{X}):=\gamma_{k}^{y}(\bm{X})-\gamma_{k^{\prime}}^{y}(\bm{X})$, we thus have
[…]
We choose $\psi(Y,\bm{W},\bm{X})$ as
$\psi(Y,\bm{W},\bm{X})=\frac{W_{k}}{\pi_{k}}(1\!{\rm l}_{\{Y\leq y\}}-\gamma^{y}_{k}(\bm{X}))-\frac{W_{k^{\prime}}}{\pi_{k^{\prime}}}(1\!{\rm l}_{\{Y\leq y\}}-\gamma^{y}_{k^{\prime}}(\bm{X}))+\gamma_{k,k^{\prime}}^{y}(\bm{X})-\Delta^{DTE}_{k,k^{\prime}}(y).$ (A.21)
Notice that $\psi$ satisfies (A.20) and that $\psi$ lies in the tangent space $\mathcal{T}$ given in (A.19). Since $\psi$ lies in the tangent space, the variance bound is given by the expected square of $\psi$:
$\mathbb{E}[\psi(Y,\bm{W},\bm{X})^{2}]=$ $\mathbb{E}\bigg{[}\Big{(}\frac{W_{k}}{\pi_{k}}(1\!{\rm l}_{\{Y\leq y\}}-\gamma^{y}_{k}(\bm{X}))-\frac{W_{k^{\prime}}}{\pi_{k^{\prime}}}(1\!{\rm l}_{\{Y\leq y\}}-\gamma^{y}_{k^{\prime}}(\bm{X}))+\gamma_{k,k^{\prime}}^{y}(\bm{X})-\Delta^{DTE}_{k,k^{\prime}}(y)\Big{)}^{2}\bigg{]}.$
When $G_{Y(k)|\bm{X}}$ is correctly specified for $k,k^{\prime}\in\mathcal{K}$, the asymptotic variance of the regression-adjusted DTE estimator coincides with the efficiency bound in the above equation.
Furthermore, the semiparametric efficiency of the DTE estimator naturally extends to QTE and PTE estimators under suitable regularity conditions. Since both QTE and PTE can be obtained through Hadamard differentiable transformations of the DTE estimator, they achieve their respective semiparametric efficiency bounds as shown in Theorem 25.47 of van der Vaart (2000).
@@ Appendix B Additional Simulation Results
(DGP1, continuous outcome, $\pi_{1}=0.3$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
(DGP2, continuous outcome, $\pi_{1}=0.3$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
(DGP2, continuous outcome, $\pi_{1}=0.5$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
(DGP3, discrete outcome, $\pi_{1}=0.3$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
(DGP4, discrete outcome, $\pi_{1}=0.3$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
@@
(DGP4, discrete outcome, $\pi_{1}=0.5$) Refer to caption Note: Bias, RMSE, 95% CI length and coverage probability calculated over 10,000 simulations. Each boxplot represents the distribution across locations $y$ for a specific sample size.
B.1 Bootstrap confidence intervals
We evaluate the finite sample performance of empirical bootstrap confidence intervals outlined in Algorithm 1 in Section 3.2 by examining their average length and coverage probability across 1,000 simulations. Figure 11 illustrates these results for DGP 1 (continuous outcome) with $\pi_{1}=0.5$. Standard errors are computed using two methods: bootstrap standard deviations (SD) and bootstrap interquartile ranges (IQR). The average lengths of the confidence intervals for both methods are comparable to those obtained using analytic standard errors shown in Figure 1. Coverage probabilities remain close to 0.95, though slight under-coverage (around 0.93) is observed at certain locations when $n\in\{300,500\}$.
Figure 11: Bootstrap confidence intervals
(DGP1, continuous outcome, $\pi_{1}=0.5$) Refer to caption Note: Average length and coverage probability of 95% bootstrap confidence intervals (500 bootstrap repetitions, 1,000 simulations). Each boxplot shows the distribution across locations $y$ for a given sample size. The first row uses bootstrap standard deviations (SD) for standard errors; the second row uses bootstrap interquartile ranges (IQR) rescaled by the normal distribution.
Figure 12 displays the results for DGP3 (discrete outcome) with $\pi_{1}=0.5$. Again, the average lengths of the confidence intervals for both bootstrap methods are comparable to those obtained using analytic standard errors shown in Figure 2. Coverage probabilities remain close to 0.95 in this case as well.
Figure 12: Bootstrap confidence intervals
(DGP3, discrete outcome, $\pi_{1}=0.5$) Refer to caption Note: Average length and coverage probability of 95% bootstrap confidence intervals (500 bootstrap repetitions, 1,000 simulations). Each boxplot shows the distribution across locations $y$ for a given sample size. The first row uses bootstrap standard deviations (SD) for standard errors; the second row uses bootstrap interquartile ranges (IQR) rescaled by the normal distribution.
B.2 Covariate transformations
In the simulation study, we examine linear regression and logistic regression models without covariate transformations, i.e., with $T(X)=X$. In this section, we explore the impact of polynomial transformations on the results. Figures 13 and 14 present the simulation outcomes for DGP1 under $\pi=0.5$, where regressors are transformed using polynomials of degree 2 and 3, respectively.
Figure 13: Performance metrics of simple and regression-adjusted DTE estimators: polynomial covariate transformation (degree 2)
(DGP1, continuous outcome, $\pi_{1}=0.5$) Refer to caption Note: Bias, RMSE, 95% bootstrap CI length and coverage probability calculated over 1,000 simulations. Standard errors are calculated as bootstrap standard deviation with 500 repetitions. Each boxplot represents the distribution across locations $y$ for a specific sample size. Regression adjustment is based on linear regression and logit model with covariates transformed with polynomial of degree 2.
For both linear regression and logit model, incorporating polynomial transformations produces results that closely resemble those obtained without transformations. Specifically, for linear adjustment, a polynomial of degree 2 achieves a 12–22% reduction in RMSE, while a polynomial of degree 3 yields a 9–24% reduction across all sample sizes. As for logit model, a polynomial of degree 2 achieves a 11–28% reduction in RMSE, while a polynomial of degree 3 yields a 10-29% reduction across all sample sizes.
Figure 14: Performance metrics of simple and regression-adjusted DTE estimators: polynomial covariate transformation (degree 3)
(DGP1, continuous outcome, $\pi_{1}=0.5$) Refer to caption Note: Bias, RMSE, 95% bootstrap CI length and coverage probability calculated over 1,000 simulations. Standard errors are calculated as bootstrap standard deviation with 500 repetitions. Each boxplot represents the distribution across locations $y$ for a specific sample size. Regression adjustment is based on linear regression and logit model with covariates transformed with polynomial of degree 3.